### v1

# Wave-induced drift in third-order deep–water theory

Abstract
The goal of this work is to investigate particle motions beneath unidirectional, deep-water waves up to the third-order in nonlinearity. A particular focus is on the approximation known as Stokes drift, and how it relates to the particle kinematics as computed directly from the particle trajectory mapping. The reduced Hamiltonian formulation of Zakharov and Krasitskii serves as a convenient tool to separate the effects of weak nonlinearity, in particular the appearance of bound harmonics and the mutual corrections to the wave frequencies. By numerical integration of the particle trajectory mappings we are able to compute motions and resulting drift for sea-states with one, two and several harmonics. Overall, we find the linear theory yields the largest forward drift close to the surface, followed by second and third order. At greater depth this trend reverses, with linear theory yielding the smallest forward drift. Incorporating difference harmonic terms into the classical Stokes drift formulation is found to yield improved agreement with the drift obtained from nonlinear wave theories, particularly at greater depth.
1 Introduction
Waves on the surface of water are a continuing source of fascination, as well as posing an important scientific and engineering challenge, with impacts on such diverse areas as offshore renewables and marine ecology. While we can observe the myriad forms of waves on the water surface, these shapes are only manifestations of what is going on below the surface: a transport of energy and, to a lesser extent, the movement of water particles themselves. Leonardo da Vinci already observed (Mossa, 2021) that the wave outruns the water, and early mathematical analyses of wave kinematics seemed to confirm this idea. Work in the early 19th century by Green showed that particle paths in linear deep water waves are approximately circular – in seeming agreement with the exact, deep-water solution derived by Gerstner – and the corresponding finite depth particle paths were shown to be approximately elliptical by Airy (Craik, 2004).
Nonlinear wave theories began to be explored by Russell, Kelland, and most prominently, Stokes, starting in earnest in the 1840s. It was the work of Stokes that first suggested that particle trajectories in periodic waves in deep water might not be closed, and gave rise to the idea of wave-induced motion now known as Stokes drift. In the intervening years the kinematics of nonlinear waves has received significant attention, and the concept of wave-induced drift has taken a prominent role in our understanding of transport in the ocean. Its consequences are still being explored in the movement of floating microplastics (van Sebille et al., 2020) and water-borne bacteria (Ge et al., 2012), and we are coming to develop a fuller picture of particle trajectories in nonlinear wave equations (Carter et al., 2020; Curtis et al., 2018; Ige & Kalisch, 2024).
Nevertheless, much remains unknown from both an experimental and theoretical perspective. Experiments are notoriously difficult, owing to the challenge of comparing a closed flume with the situation in the (open) ocean. Recent years have seen significant progress, from flume experiments on monochromatic wave trains (Umeyama, 2012) to more recent investigations of particle kinematics for groups and complex flows (Bjørnestad et al., 2021; van den Bremer et al., 2019).
The aim of this work is to explore the theoretical particle trajectories – and drift – associated with nonlinear potential flow theories in infinite water depth. While there are recent mathematical results for the full, nonlinear problem showing that no closed particle paths exist in steady, periodic wave trains (Constantin, 2006; Henry, 2006; Okamoto & Shoji, 2012), these results are not quantitative and are difficult to generalise to nonsteady problems. We will take a ‘weakly nonlinear’ approach, and our starting point will be the Stokes drift formulation which arises in linear (1st order) water wave theory. While a rigorous proof that linear, steady, periodic water waves exhibit a forward drift in the direction of wave propagation (Constantin & Villari, 2008; Constantin et al., 2008) was recently found, the approximate results due to Stokes form the basis of most practical implementations of drift in forecasting codes.
Stokes’ original results for a single wave train were generalised to a spectrum of waves by Kenyon (1969), and the resulting formulation has been extensively explored for different spectral shapes (Webb & Fox-Kemper, 2011), for its sensitivity to the high-frequency spectral tail (Lenain & Pizzo, 2020), and new profiles have been developed for practical implementation (Breivik et al., 2016). Many more references can be found in the recent review by van den Bremer & Breivik (2018).
As the Stokes drift formulation comes about through a simplification of the linear water-wave theory, the differences between it and the linear theory are of interest. This also raises a natural question: what if we progress from linear theory to a second or third–order theory? We expect the higher-order theories to be more accurate, and to include important effects for steeper waves – such as the appearance of additional harmonics – which may have a bearing on comparisons with the Stokes drift. Indeed, these theories may suggest improvements to the Stokes drift derived from linear theory. However, with the exception of recent work of van den Bremer and Taylor (2016) focusing on wave groups it is difficult to find such comparisons in the literature.
For a successful comparison we need some way to compute the fluid velocity field , which drives the particle kinematics through the particle trajectory mapping We need a methodology that can deal with several Fourier modes, to simulate irregular seas, and which is accurate to a desired order of nonlinearity. Several options exist, including the use of traditional perturbation theory to obtain a solution to the water wave problem, which has been extended to third-order by Madsen & Fuhrman (2012). An alternative – which we employ – is to use the equivalent Hamiltonian expansion, which is due to Zakharov (1968) and for which the third-order solution has been explored by Zhang & Chen (1999) and Gao et al. (2021).
The remainder of this paper is structured as follows: in Section 2 we introduce the water wave problem and corresponding Hamiltonian. We give more details about the expansion that allows us to extract linear, second, and third-order terms in Section 2.1, discuss the neglect of amplitude evolution for our particle kinematics calculations in Section 2.2, and finally show how the free surface and potential are obtained to each order in Section 2.3. In Section 3 we consider monochromatic waves, which is the natural setting to introduce Stokes drift, and look at the consequences of higher-order theories on particle trajectories. Section 4 is devoted to bichromatic waves. These are the simplest type of non-steady wave-field, but because only two harmonics are involved the linear superposition is always periodic. Finally, Section 5 considers wave groups and random waves involving multiple lowest-order harmonics. Conclusions are presented in Section 6. Appendix A adds some comments on the non-uniqueness of the monochromatic Stokes wave solution at third-order, while Appendix B gives expressions for some coefficients appearing in Section 2.3.
2 The water wave problem and Hamiltonian formulation
The inviscid, irrotational water wave problem, written in terms of a fluid velocity potential and free surface is
The horizontal () gradient is denoted for clarity. The advantage of this formulation is that all nonlinearities have been moved to the boundary conditions. We also note that, provided the water is deeper than about half a typical wavelength, the still-water depth can be taken to be infinite. This procedure, using
(1d’) |
in place of (1d) simplifies calculations considerably. Despite the restriction to irrotational flow and the neglect of compressibility and viscosity, these equations contain a huge range of interesting physics, including many types of wave phenomena.
During the late 1960s, efforts to find a Hamiltonian formulation of the water wave problem (1a)–(1d) came to fruition in the independent work of Zakharov (1968), Miles (1977) and Broer (1974), who found the Hamiltonian
which represents the total energy of the fluid. The domain of integration in can in practice depend on the specifics of the problem, and the lower boundary of the -integral can be chosen as if (1d’) is used in place of (1d).
The canonical variables for the Hamiltonian formulation are defined at the free surface. They are and and the canonical equations are thus
These equations are equivalent to the surface boundary conditions (1b), (1c), and details of the calculation can be understood most easily in Broer (1974). Additional background can also be found in the recent review by Stuhlmeier (2024).
2.1 Asymptotic analysis and the water wave problem
The water wave problem, whether written as a system of partial differential equations and boundary conditions or using the Hamiltonian reformulation remains mathematically very complex. While the mathematical existence of periodic and solitary wave solutions has been established in ever more general settings (for example, the books by Constantin (2012) or Lannes (2013) offer a relatively recent overview of mathematical progress), much of what is known about qualitative and quantitative properties of ocean waves rests on an analysis of simplified equations.
These simplified equations are usually obtained through so-called asymptotic analysis or perturbation theory – a procedure of nondimensionalisation and scaling the equations, followed by the identification of small dimensionless parameters, and an asymptotic expansion of the variables in terms of the small parameters selected. In the case of water waves the most common small parameter is the wave steepness written either as the wave amplitude divided by the wavelength or multiplied by the wavenumber. A very thorough discussion of asymptotic analysis as applied to water waves can be found in the textbook by Johnson (1997).
Instead of expanding the water wave equations (1a)–(1d) in this manner (see Madsen & Fuhrman (2012) for a detailed, relevant account), it is also possible to expand the Hamiltonian directly. This procedure was carried out by Zakharov (1968) and, with more detail, Krasitskii (1994), and it has the advantage of enforcing structure that might otherwise be lost, such as that resulting equations conserve energy.
The first step in this procedure is to write the canonical equations in terms of the Fourier transformed variables and as
with denoting the complex conjugate. Then one defines a new pair of canonical variables and related to and via Fourier transforms as
where
is the frequency (in rad/s), and is the constant acceleration of gravity (taken to be 9.8 m/s2 in computations). As we will be interested in deep-water waves (, we note that the dispersion relation simplifies to
We will use this in all the computations that follow.
The subsequent steps in the expansion are somewhat lengthy, and are given by Krasitskii (1994). At this stage it is important to emphasise only one key point: at lowest order the problem is linear, so the superposition principle holds. This means that the Fourier amplitudes – which correspond essentially to the magnitudes of the terms introduced in (4)–(5) – do not change with time. If gravity is the sole restoring force, the surprising fact is that the Fourier amplitudes do not change with time at the next order, either. Only at third-order in nonlinearity, where resonance between quartets of gravity waves becomes possible, do we obtain an evolution equation. This equation gives the slow evolution of the complex amplitude
and is known as the Zakharov equation, with a kernel term given in (Krasitskii, 1994, (3.5)). We have introduced subscript notation to denote wavenumbers, so that, for example, the delta-function is a shorthand for In principle can be either a scalar or a vector wavenumber, although our computations will deal only with the scalar case for simplicity.
If is known then the canonical variable can be recovered via
where we have written terms up to third order only. This expression is given in (Krasitskii, 1994, (2.17)), as are the relevant interaction kernels . Finally, once is known it is possible to recover and up to the desired order by means of (4)–(5).
2.2 Constant magnitude approximation
The procedure outlined above would seem to suggest that we first need to solve the complicated integro-differential Zakharov equation (8). In fact, for certain situations of interest we can circumvent this step. The key insight is that the evolution described by the Zakharov equation occurs on the slow timescale , where is a typical wave steepness and is a typical wave period (e.g. s in the ocean). Therefore, for processes taking place on time-scales faster than 100 wave periods, it may be appropriate to consider the complex amplitudes to be constant.
This is indeed the approach we will adopt. There is, however, one important addition: for certain wavenumber combinations the Zakharov equation (8) can be solved explicitly. In particular, for all symmetric quartets which satisfy or the equation reduces to a system of one or two ordinary differential equations with constant-amplitude solutions, as shown in (Mei et al., 2018, Ch. 14). Such solutions represent nonlinear corrections to the frequency of the waves, and were first found by Stokes for a single wave mode. For arbitrarily many modes, or a continuous spectrum, it is possible to write these corrections in a compact form (Stuhlmeier & Stiassnie, 2019).
Incorporating these corrections the complex amplitudes can be written in an approximation up to third-order as
where the small, quadratic terms in the argument reduce to , the well-known Stokes’ correction, when only a single wave-mode is present. This form now contains all the mutual dispersion corrections entailed by the cubically nonlinear water wave theory. It has been demonstrated that taking these corrections into account allows one to predict the short-time propagation of laboratory waves with high accuracy without the need to solve for the evolution of the slow-time variables (Stuhlmeier & Stiassnie, 2021; Galvagno et al., 2021), showing that the constant magnitude approximation is well-justified for short times. Of course, under specific circumstances – for example with a sufficiently narrow spectrum, or close to modulation instability – we may have significant energy transfer (Andrade & Stuhlmeier, 2023). However, given our focus on particle paths (with a characteristic time on the order of one period) the amplitude evolution is negligible (see also the discussion in Stuhlmeier & Stiassnie (2019) and Appendix A therein).
2.3 Recovery of the third-order solution
The Hamiltonian formulation makes clear how the free surface and potential can be recovered from the canonical variables and up to a given order. In practice, discrete wavenumbers are considered, which means that the integrals in (9) can be simplified by the substitution
Here is unknown in principle, and in practice comes from a constant magnitude approximation to (8) as discussed in Section 2.2.
Because the free surface (4) is recovered directly from substitution of (9), it is clear that we can distinguish linear, quadratic and cubic terms in based on the respective terms in (9). The same is not true of , as we shall see below. Taking the inverse Fourier transform yields
Here we have substituted with amplitude and phase and employ the compact notation Clearly this yields the expected sinusoidal free surface elevation upon defining the amplitude
The nonlinear corrected frequency is given by
and for unidirectional waves in infinite depth – which will be our focus – we can simplify the kernels to
The quadratic components of the free surface elevation can be obtained in analogous manner, with slightly more in the way of algebra required to resolve all delta-functions. We find
The shorthand notation is used to denote
We can write the cubic components of the free surface elevation as
The velocity potential at the free surface, , is obtained from (5) via an identical procedure.
The corrections contained in the second-order and third-order free surface elevation essentially change the wave shape. The resulting harmonics of the form behave as waves with wavenumber and frequency but because these sums and differences do not satisfy the dispersion relation they are referred to as “bound waves”, in contrast to the free waves of (11). The well-known impact of these bound waves on sharpening the crests and flattening the troughs is shown in Figure 1 for a bichromatic sea, together with the effect of the third-order frequency correction (12) on the wave phases.
2.3.1 Recovery of the bulk potential
The reformulation of the problem in surface variables is very convenient from many perspectives, but is something of a disadvantage when we wish to recover information about the bulk fluid below the free surface. To obtain this we start with the general solution of the Laplace equation which decays for infinite depth
the exact analogue of (Krasitskii, 1994, (4.1)).
We relate the bulk potential and the surface trace to one another in terms of an expansion. Taking the problem in infinite depth, we need to solve
which we do by assuming the free surface elevation is small (), and then inserting this expansion into the boundary condition (17), so that
This gives rise to a hierarchy of problems in physical space, whose successive solution yields
with
With and known from the formulation in (4)–(5) we can now recover the bulk potential at each order desired. This procedure is also detailed in (Janssen, 2004, below (4.7)), (Krasitskii, 1994, p.15) and (Zakharov, 1968, (1.8)), although care should be taken in the symmetrisation of (20).
The linear contribution consists simply of the first term in (19)
The quadratic contribution is
where we need to employ the linear and quadratic parts of and (and products thereof) in the formulation. The coefficients are given in Appendix B. Up to second order, for unidirectional waves in deep water the potential reduces to the simple expression
This follows also from Dalzell (1999).
The cubic contribution to the bulk potential is
The coefficients are given in Appendix B. Unfortunately efforts to find a compact simplification the third order have failed except in the case of a single monochromatic wave, where the analytical expression
3 Particle kinematics for monochromatic waves
We have developed above a general formulation of the third-order water wave problem. However, this formulation in terms of a free surface elevation and a velocity potential needs to be connected to the kinematics of fluid particles. The most transparent starting point is with monochromatic waves – i.e. initial conditions for which the linear theory is described by a single Fourier mode Of course, the lengthy formulation of Section 2.3 is not strictly necessary, since monochromatic solutions to third-order can be readily found in the literature, e.g. Wehausen & Laitone (1960). Consequently, we begin by reviewing some familiar ideas.
3.1 Linear theory
Linear deep-water theory poses the water wave problem on the half-plane below Solving the Laplace equation then yields the well-known solution in deep water
where and
is the linear dispersion relation. Here and
The particle trajectory mapping is the system of ordinary differential equations
It is worth noting that this system has only a tenuous connection to the linear (or weakly nonlinear) theory, but rather represents the link between Lagrangian and Eulerian descriptions of the fluid motion. For example, the idea that the free surface is always composed of the same fluid particles is only nearly satisfied (under the assumption of waves of small amplitude), so that the particle trajectory mapping only nearly describes the dynamics of particles on the surface (or any streamline).
Given that we are dealing with linear theory, it seems simple enough to solve the system (26). It is an unwelcome surprise that inserting the linear potential yields a nonlinear system of differential equations
While this system cannot be solved analytically it can be shown rigorously that its particle trajectories are not closed (Constantin & Villari, 2008).
In lieu of an analytical treatment, two alternatives remain: approximation or numerical solution. The former approach is found in most textbooks on water waves, and begins with a Taylor expansion of the fluid velocity field as
and similarly for Keeping only the lowest order terms (i) in the Taylor expansion yields the explicit system
with which yields the circular particle trajectories first found by Green in 1839 (Craik, 2004). Inserting the solution so obtained into the particle trajectory ODEs with terms in (29) retained yields another explicit system
Integration of this system yields the familiar result that particles move forward with a Stokes drift velocity
This Stokes drift is nominally a second order – and therefore nonlinear – quantity in the small wave amplitude , but is derived from linear theory (24). This important reminder can also be found in many textbooks, such as Dean & Dalrymple (1991).
Despite the familiarity of the foregoing, a few remarks are in order: the system (32)–(33) which yields the Stokes drift is an approximation of (27)–(28), under the assumption that each particle remains close to its initial position . The result of integrating both systems – (32)–(33) analytically (dashed curves) and (27)–(28) numerically (solid curves) – is shown in Figure 2. Due to the explicit time dependence in the former, the time needed for each particle motion to return to its initial vertical position is exactly equal to the Eulerian period Moreover, after the passage of one Eulerian period each particle has moved an amount in the direction of wave propagation. This drift is indicated by the dashed black curves, which connect the end-points of the dashed trajectories in Figure 2.
By contrast, the time needed for a particle to return to its initial vertical position in the exact, linear theory – which we call the Lagrangian period – is always larger than the Eulerian period. The Lagrangian period is largest at the top of the fluid domain and tends towards the Eulerian period as , as shown also by Constantin & Villari (2008). Moreover, the forward drift of a particle depends on the initial condition, as shown in the black curves connecting the end points of the linear particle trajectories in Figure 2. This drift is sometimes larger and sometimes smaller than the Stokes drift, though its computation using must distinguish between and in the denominator and the corresponding horizontal displacement after a Lagrangian period (for (27)–(28)) or an Eulerian period (for (32)–(33)). Averaging over initial positions covering one wavelength, we find that integration of (27)–(28) yields a larger forward drift than the 2nd order Taylor series approximation (34), as summarised in Figure 3 for several different slopes.
3.2 Nonlinear theory
It is well-known that the deep water potential at third order is formally identical to the first order potential given in (24) (though see Appendix A). Thus, while higher harmonic terms appear in the free surface elevation, the only change to the potential at third order is the nonlinear frequency correction, which means that the characteristic coordinate is now for
In particular, this means that the speed at which each particle moves is faster by an amount of order which impacts the particle trajectories as shown in Figure 4.
For monochromatic waves the procedure used to derive (34) can be applied without alteration to yield the nonlinear corrected Stokes’ drift velocity
Because it is clear that the drift velocity (36) is generally slower than (34) – this despite the fact that nonlinear waves travel faster than their linear counterparts. Moreover, for fixed steepness this correction is more significant for longer waves, i.e. smaller wavenumbers This counterintuitive result – that despite nonlinear waves travelling faster their drift velocity is slower – is due to the fact that the particle trajectories themselves “shrink” with increasing . To lowest order in the Taylor series expansion, where deep-water trajectories are approximately circular, their radius becomes
For the cases considered in Figure 3 the introduction of the frequency correction when leads to a 2 % reduction in the approximate Stokes drift ((36) vs. (34)) and a 5 % reduction in the averaged Lagrangian drift obtained from integrating the particle trajectory mapping with the third-order potential inserted. We will encounter this unexpected consequence of nonlinear theory again in our discussion of more complex wave fields below.
4 Bichromatic waves
In the previous section we considered monochromatic waves. The term “monochromatic” refers to the fact that a single mode is assumed in the lowest order problem. As we have seen, higher harmonics are excited at each order of the perturbation expansion, so that the spectrum does not, in fact, remain monochromatic. Far from being a purely mathematical phenomenon, this harmonic generation is observed every time a wavemaker is excited in a flume, and remains an active area of experimental research (Khait & Shemer, 2019). Consequently, when we speak of bichromatic waves, this means that two modes and are assumed in the lowest order problem. The second order solution then contains sum and difference terms of any two of these, while the third order contains sum and difference terms of any three, with an attendant increase in complexity. Up to second order the solution is contained in the work of Dalzell (1999), while the third order solution for bichromatic waves appears first in the work of Madsen & Fuhman (2006). We make use of the equivalent formulation of Section 2.3.
4.1 Linear theory
We start with the simplest scenario, where we retain only linear terms in the bichromatic velocity potential
To understand what this means for particle kinematics, we are again confronted with two options: solve the particle trajectory mapping numerically or approximate it in the same fashion as the monochromatic Stokes’ drift in Section 3.
Exactly as for monochromatic waves, if the lowest order terms in a Taylor series expansion of (37) are retained, the resulting system with explicit time dependence
is trivially integrated. Substituting the solution into the Taylor series expansion (cf. (29)) leads to a somewhat lengthy term; while the products of the first order solution and the velocity field produce terms involving the time dependence of the equations is, of course, explicit. Upon integration the secular terms lead immediately to the drift velocity
which is simply the sum of the Stokes’ drifts for modes and over a time t. This simple idea, that the Stokes drift of a linear sum of waves is the sum of the Stokes drifts of the waves underlies the successful application of Stokes drift to spectra as pioneered by Kenyon (1969).
As we have for monochromatic waves, it is worth comparing the Stokes drift formulation (40)with the actual particle trajectory mapping. Such a comparison is shown in Figure 5, for different snapshots of one bichromatic wave-train. Although it is mathematically elementary, it is worth some digression to explain this figure. Bichromatic wave-trains look quite different in space and in time, so we choose to visualise the free surface as a function of time at different spatial locations. Two radian frequencies rad/s and rad/s are considered, such that their superposition has a (Eulerian, temporal) period of s. (With m/s2 the bichromatic wave train also has a spatial period of approximately 986.2 m.) Several snapshots of the free-surface time-series over one such period are shown in Figure 5 panels (b)–(d), with the envelope indicated by the pair of dashed, black curves.
Panel (a) of Figure 5 shows the trajectories of particles starting at and various depths ( and -4.8 m) over one temporal period. Exactly as in Figure 2, the solid curves are obtained by integrating the particle trajectory mapping with the bichromatic velocity field obtained from (37), while the dashed curves shown an analytical solution obtained from retaining the 2nd order terms in the Taylor series expansion of the same equations. Consequently, the forward movement of a particle along the dashed trajectory is given simply by (40), shown as a black, dashed line in Panel (a), while the forward movement of a particle computed from numerical integration is shown as a black solid line
Panels (b)–(d) show the free surface elevation at three spatial locations, and the panels below each show the corresponding trajectories (panel (e) shows an enlarged version of the blue trajectories in panel (a) with initial position ). Note that, while the free surface envelope in panels (b) and (d) may look very similar, the trajectory in (e) begins at a crest while the trajectory in (g) begins near a trough. This difference is clearly visible in the trajectories themselves. Trajectory (f), on the other hand, begins near a local minimum (in space) close to the mean water level
As for monochromatic waves, we can compare the computed Stokes drift (40) with the values obtained directly from integration of the particle trajectory ODEs. For the example considered in Figure 5 we observe (as in Section 3) that the Lagrangian period is always larger than the Eulerian period. Averaging the drift velocity over one spatial period yields Figure 6, where dashed curves represent the Stokes drift velocity and solid curves the computed averaged Lagrangian drift velocity. Bichromatic wave-trains with rad/s and rad/s and and 0.1 are considered. The same qualitative features as in Figure 3 are observed: linear theory predicts a larger forward drift than the Stokes’ drift approximation. This is also the case when the bichromatic wave modes have different slopes.
4.2 Nonlinear theory
4.2.1 Second order
If linear bichromatic waves are very simple, the second order expression obtained from (22) is not much more complicated:
where and the sum harmonic vanishes for unidirectional deep water waves (the same is not true of the free surface elevation, whose prominent sum harmonics can be seen in Figure 1).
The second order solution (41) lends itself to a calculation of the Stokes drift as outlined in Section 3.1: calculate , Taylor expand about a point , keep only the constant terms in the Taylor expansion and calculate approximate trajectories , subsequently substitute these trajectories into the linear terms in the Taylor expansion and integrate the resulting particle trajectory equations in time. The result of this procedure is
which consists of the term (40) from the linear theory as well as a new term (denoted (II)) that arises from the difference harmonics (recall that we assume ).
At first sight the term (II) may seem insignificantly small, however there are regimes where it becomes important in a description of particle drift. It is simplest to analyse this by fixing 1/m and varying for At the top of the fluid domain we observe that the contribution (II) grows monotonically with depicted as the dashed, yellow curve in Figure 7(a). A consequence of this is that there is minimum of the total 2nd order Stokes drift (42) for a particular value of For we find that term (II) has a single maximum and subsequently decays, as shown for two depths in Figure 7(b)–(c). As we descend into the fluid, the maximum of term (II) moves towards smaller values of
While overall particle motion decreases the deeper we descend into the fluid, there are regions where the 2nd order contribution to Stokes’ drift has a sizable effect (see Figure 7(b)) and in fact dominates the 1st order contribution (I), as shown around in Figure 7(c). This is in line with the importance of difference-harmonic terms in generating dynamic pressures that reach the ocean bed (Longuet-Higgins, 1950).
To further understand the importance of the second order Stokes’ drift correction, we compare the Stokes’ drift formulation with the particle drift obtained from integration of the particle trajectory mapping. Panels (d)–(f) of Figure 7 show the average horizontal drift obtained by integrating using the 1st order velocity field (37) or the 2nd order velocity field (41). The averaging is performed over initial conditions spanning one wavelength of the bichromatic envelope, and both modes have identical steepness .
Panel (d) depicts the drift at for a bichromatic wave train with 1/m and 1/m. The discrepancy between the linear Stokes drift (I) and 1st order solution has already been noted in Figure 6, and is seen again in comparing the solid and dashed blue lines. It is therefore rather surprising that incorporating the second order Stokes drift (42) leads to much better agreement with the exact, numerical solution, as shown by the dashed and solid red curves.
In panels (e) and (f) of Figure 7 we consider a bichromatic wave train with 1/m and 1/m, close to the maximum of the yellow curve (II) shown in panels (b) and (c) at depths of m and -5 m, respectively. At such depths the overall particle motion is much smaller, but for a range of wavenumbers the 2nd order component dominates the Stokes’ drift. Integrating the particle trajectory mapping bears this out: the blue curves showing the expression (I) overlap with the solutions of the particle trajectory mapping with 1st order velocity field, as in Figure 6. Significantly, the red lines showing the analytical expression (I)+(II) given in (42) also overlap with the respective 2nd order solutions. It can be checked that for where contribution (II) is insignificant, all four lines overlap.
4.2.2 Difference harmonics and particle drift
Why do we observe this reversal, where linear theory predicts a larger drift near the surface, but second order theory predicts a larger drift at depth? The answer lies in the structure of the difference harmonic terms, and as a prelude to our work on wave groups it is expedient to investigate their role in bichromatic seas.
The first consideration should be of the velocity field which drives the particle trajectory mapping. We consider the waves studied in Figure 7 with and 1/m, and calculate horizontal velocities directly from (41). When only the first order contributions are retained we find the familiar pattern of forward velocity below the crests and backward velocity below the troughs, with an exponential fall-off in depth as depicted in the top panel of Figure 8.
If instead we focus only on the second-order contributions , the situation looks quite different. The exponential decay in velocities is slower because is larger than either or for and we find a negative horizontal velocity under the centre of the bichromatic wave train, as shown in the bottom panel of Figure 8. The full second-order picture is a sum of the first and second order terms, shown in the middle panel of the same figure, and it is clear that the first order dominates near the surface while the second-order difference harmonic dominates at depth. This can be observed clearly by considering the zero contour lines of the horizontal velocity, shown in black for each case.
The generic situation for particle paths in a bichromatic wave train is thus as follows: close to the surface the first order contributions dominate, giving the curlicue particle paths observed in Figure 5 and in Figure 9(b). The effect of difference harmonics is to retard slightly the forward motion of the particles, due to the slow moving (in deep water the group velocity is half the phase velocity) region of backward (negative) velocity under the highest crests observed in Figure 8. This retardation from second-order difference terms gives rise to the larger drift in the linear theory, observed in Figure 9(c).
Deeper into the fluid column the effect of the first order free-modes becomes negligible. The fluid motion is dominated by the difference harmonic term , and the particle paths behave essentially as those of a monochromatic wave with that wavenumber. This is clearly illustrated in Figure 9(d), where we see that the modes and which make up the first-order theory are only a small perturbation on the displacement due to the difference harmonic term.
It is worth pointing out the existence of a region beneath the bichromatic wave train where the particle displacement is largely against the direction of wave propagation, as seen between in Figure 9(e), and indicated with yellow circles. This should not be surprising, since it simply reflects the effective motion of the difference harmonic term as shown in panel (d). Despite the existence of this ‘return flow’ – which we shall encounter again in our discussion of wave groups – the net particle motion remains forward at every depth in the fluid, as shown in Figure 7 and Figure 10 below. Indeed, the relatively large displacement associated with difference harmonic terms at depth gives rise to the larger drift observed for the second-order solution.
4.2.3 Third order
A confluence of the effects of dispersion correction and harmonic generation occurs for the first time at third order. For the frequency correction appearing there we have the notably asymmetric equations
for which were first found by Longuet-Higgins & Phillips (1962). Note the presence of the Stokes correction found in (35), and the fact that the longer wave () has a much more significant effect on the shorter wave () than vice versa.
The higher harmonics (or bound waves) have the well-known effect of sharpening wave crests and flattening wave troughs, as shown in Figure 1. To first order (dark blue curve) this figure shows the same bichromatic wave-train treated in Section 4.1, albeit with the slope of each mode increased to for ease of visualisation. Note that the 1st and 2nd order solutions (shown in dark and light blue) employ linear dispersion. The 3rd order solution (shown in red) includes the mutual frequency correction (43)–(44), and so has a different period than – and in fact is seen to move out of phase with – the lower order solutions.
While a compact form of the third-order bichromatic potential akin to (41) is not available, it is relatively simple to simulate the exact, third-order theory on a computer using the results of Section 2.3. This allows us to compare the computed forward drift for the same bichromatic wave train using the 1st order velocity field, the 2nd order velocity field, or the 3rd order velocity field, as well as the various formulations of Stokes drift. Such a comparison is shown in Figure 10, which considers the depth-dependence of the wave trains shown in Figure 7. Because the particle drift depends on the initial horizontal position – i.e. at which point in the phase the wave train is encountered – this figure is calculated by averaging over one spatial period at each depth.
Close to the surface we observe the largest drift in 1st order theory, followed by 2nd order and 3rd order. Incorporating difference harmonic terms (II) into the Stokes drift leads to a closer agreement with the exact second and third order drift as shown in the left panel of Figure 10. With greater depth, and for nearby wavenumbers, we have a distinct reversal in the ordering: 1st order theory, whether exact or using the Stokes drift approximation, yields the smallest forward drift. The 2nd order theory gives the largest drift, with the 3rd order slightly smaller. Once again, incorporating difference harmonic terms into the Stokes drift provides a distinct improvement in such cases over depth.
5 Multichromatic waves
When more than two Fourier modes are necessary to describe the linear problem, we call such a configuration multichromatic. The second order contains sums and differences of any two modes, while the third order contains sums and differences of any three. This means that significant algebra can be involved in calculating higher-order solutions.
5.1 Focused linear wave groups
In treating waves with an increasing number of linear harmonic constituents we also have a plethora of parameters that can be tuned. One case of interest, from both an experimental and theoretical standpoint is that of a (focused) wave group. Linear focusing is quite straightforward: if our free surface is a superposition of sinusoids as per (24), then writing
readily gives a wave train that focuses – i.e. all components come into phase – at . Pulse-like wave groups can be created by suitable tuning of the amplitude spectrum One possible choice is a Gaussian amplitude spectrum of the form
An example of a wave train with such an amplitude spectrum is shown in Figure 11(a), where we have chosen equidistant modes between and 1/m, , and The focusing location is chosen to be , and the focus occurs at a crest.
In order to more clearly visualise the behaviour of the particles the curves in Figure 11 have been coloured with a time-varying gradient. Comparing panels (a) and (b) we observe that a particle at at time initially undergoes very little motion. Panel (c) shows that its horizontal position is effectively constant at until nearly s, as is its vertical position (not shown). As the group passes the particle moves with the wave motion, before settling at a new horizontal position close to m.
At greater depth the particle motion from the linear wave group is shown in Figure 11(d)–(e) where m. Here the particle seems very nearly to return to its initial position along a spiraling trajectory shown in panel (d), with nearly vanishing forward drift, as expected from our discussion of monochromatic and bichromatic waves.
5.2 Nonlinear focused waves
The particle trajectories obtained by integrating the linear velocity field in Figure 11 can easily be recalculated using either second or third-order theory. Experience with bichromatic waves in Section 4 teaches us to expect that all theories will behave similarly near the free surface, with linear theory giving the largest forward drift followed by second and third order. Indeed this is what we observe, as shown in Figure 12(a)–(b), where the colour gradient is the same as in Figure 11.
In contrast to the bichromatic case, we now have a plethora of difference harmonic terms which have an influence further down the water column. Difference harmonics close to for the group dominate, and lead to a 25-fold increase in the displacement at depth when compared to the linear theory, as shown in Figure 12(c)–(d). We also observe the return flow associated with these trajectories, exactly as for bichromatic waves in Section 4.2.2. Interestingly, the third-order contribution, consisting of both a change in dispersion relation as well as the presence of additional harmonics, has a notable effect on the particle displacement throughout.
We can get a better idea of the depth-dependence of the particle motion from Figure 13, which shows the horizontal position of a particle initially at m after the passage of the wave group discussed above. First, second and third-order solutions are denoted by blue, red and yellow curves, respectively, and the Stokes drift (generalising (40) to 10 modes) is denoted by a purple curve with label ’S’. This figure clearly illustrates the role of higher-order theories in reducing the drift close to the surface, and in inducing a significant return flow beneath the group with depth. Indeed, while linear theory and Stokes drift are reasonably close in Figure 13(b), they cannot capture the sizable negative velocities beneath the group.
Care should be taken, however, not to overinterpret this result. As we have seen for bichromatic waves in Section 4.2.2, the return flow is a highly localised phenomenon, and does not indicate a net negative drift. Our wave groups are clearly not envelope solitary waves – whose existence in the third-order water wave problem remains open (Slunyaev & Shrira, 2013) – composed as they are of a finite number of Fourier modes. Consequently there is no approach to a limiting state of zero-flow as the wave group passes, rather the passage of one group simply heralds the appearance of another following behind it. The particle motion induced by such a succession of wave groups recovers the expected forward drift.
5.3 Random waves
Random waves are the closest mathematical idealisation to the waves we might typically find in the ocean: the phases of one Fourier component are not correlated with the next, and the resulting pattern – while occasionally exhibiting a distinct group – is highly irregular. Random phases at lowest order also means that sum and difference phases at higher orders are random, due to a neglect of slow phase-coherence effects associated with nonlinear coupling (Andrade & Stuhlmeier, 2023).
Because we are interested in comparing different computations of drift, we also derive a generalised version of the Stokes drift (42) with second order difference harmonics suitable for arbitrarily many modes. The procedure is familiar: Taylor expand the velocity field from (22) about an initial position , and keep the lowest order terms. Solve for the particle trajectories by integrating that expanded velocity field in time. Then substitute the solution thus obtained into the second order terms in the Taylor expansion, and integrate the result to find the corrected particle trajectories. We obtain
Here we recognise term (I) as simply the sum of the individual Stokes drifts, and term (II) as the generalisation of (42).
Figure 14 shows the results of computing the wave-induced particle displacement for random waves. The amplitude spectrum (45) is retained, but rather than assigning the phases of the waves so as to achieve linear focusing at a specific location, these phases are randomly generated for 200 realisations. Red curves denote first order theory, green curves second order, and blue curves third order, with the prominent, dark curves representing the average over all realisations. In each case the integration is for approximately 80 peak periods, which is long enough to clearly distinguish the particle drift.
As we observed for focused waves in Figure 13 and bichromatic waves in Figure 10, near the surface the linear theory gives the largest forward drift. While particular realisations (shown in lighter curves) may exhibit a larger drift in a nonlinear (2nd or 3rd order) theory, on average the linear theory is an overestimate of the forward drift. Likewise, as was the case for focused waves and bichromatic waves, the Stokes drift formulations represent an underestimate of the drift. Unlike specific cases identified for bichromatic waves in Figures 7 and 10 there is no discernible difference in the Stokes drift when difference terms (II) are included near the surface.
The more interesting setting for these random waves is observed at greater depths, which is shown in panel (b). With a peak wavenumber 1/m, most of the motion due to the free-wave constituents is filtered out at a depth of m. The linear theory gives a very narrow range of displacements, with the realisations in red barely showing deviation from the average, and agreeing well with the linear Stokes drift (I). Second and third order displacements are quite large in certain realisations, depending on the phase relationships involved. However, as expected, there is no systematic return flow at any depth for random waves, because such (spatially and temporally) localised flows are averaged out to yield the net forward drift which we observe in Figure 14(b). Moreover, all theories approach zero from above with depth, exhibiting the same ordering we have hitherto observed: linear theory predicting the smallest drift followed by second and third order. The correction (II) in (46) is found to give a useful improvement over using only the classical Stokes drift term (I).
6 Conclusions
The wave induced drift of particles has attracted increasing attention in recent years, particularly in connection with the transport of maritime pollution. Nevertheless, fundamental explorations of this topic are rare. Notable exceptions are the recent articles by van den Bremer & Taylor (2016) and the follow-up study by van den Bremer et al. (2019), which focused on the transport associated with wave groups to second order. Our aim has been to explore particle drift systematically using the first, second, and third order theory, and working our way from monochromatic waves to random seas. This work rests on numerical integration of the particle trajectory mappings and comparison with simplified formulas like that for Stokes drift.
We consistently observe that linear theory gives the largest forward drift close to the free surface. Adding second order effects decreases this drift somewhat when more than one mode is present. The reason is simple: where the linear waves have the largest displacement we find a region of negative velocity in the second order difference harmonics. Third order effects decrease the forward drift further through a combination of changes to wave speed – which shrinks the particle paths – and the addition of further harmonics.
Descending in the water column we find a different overall behaviour of the wave-induced drift. It is well known that the water column acts like a filter, first suppressing the highest frequencies and filtering out ever lower frequencies with increasing depth. Particle motion is typically considered negligible half a wavelength below the surface. It is this filtering effect on linear, free-wave constituents that leads to the dramatic decrease of drift in the linear theory. However, nonlinearity introduces much longer bound-waves which become dominant at depth. Beneath the centre of a wave group these can induce sizable localised return flows. However, their overall contribution is to increase the drift in the direction of wave propagation versus linear theory. Third-order contributions lead to additional bound waves, but also to dispersion corrections; the overall effect of third-order vs. second-order nonlinearity thus depends on parameters of the sea-state such as its bandwidth and amplitude spectrum.
These results for each order require integration of the particle trajectory mapping, which is inconvenient. The more common approach in practice is to use the Stokes drift formula, which approximates the linear theory via a Taylor series expansion. Comparisons show that this approximation is quite poor, particularly close to the surface, where linear theory yields a significantly larger drift. In fact, the Stokes drift formulation inadvertently lies quite close to the smaller drift predictions of the nonlinear (second or third order) theories near the surface, which may explain some of its success.
At depth the Stokes drift hews close to the linear drift, but we have seen that this may differ significantly from the drift in nonlinear wave theories. To this end, we have retained a further term in the Stokes drift formulation which bridges some of this gap, and accounts for second order difference harmonics. For settings with multiple nearby harmonics this adjusted Stokes drift formula captures the larger forward drift associated with nonlinear waves, and represents an improvement on the commonly used formulation.
Our objective has been to provide a systematic and transparent exposition of drift associated with waves in nonlinear potential flow. To this end numerous simplifications have been made. Noteworthy among these is the neglect of viscous effects at the free surface, which has been known since work of Longuet-Higgins (1953) to play an important role. We have likewise neglected directional effects, which can be expected to significantly complicate any study. Finally, and significantly, we have relied on the simplification of infinite water depth. This has allowed us to present numerous simple and compact formulations which are difficult to find in the existing literature, and contributed – we hope – to a clear presentation. Nevertheless, it is important to note that long bound harmonics may “feel the bottom” even when the free waves do not, and so the inclusion of finite depth is of central importance. This will be treated in a subsequent work.
With all of the caveats above, it is important to mention that the drift incorporated into operational models such as WAVEWATCH III is the simplest infinite-depth Stokes drift formulation (Li et al., 2016). This points to a continued need to better understand this formulation – with all the simplifications it involves – and how it may be further refined. For example, it has been pointed out recently (Lenain & Pizzo, 2020) that correct treatment of the high-frequencies is critical for estimating the surface drift. The high-frequency components of the wave spectrum undergo the largest correction due to third-order effects (Stuhlmeier & Stiassnie, 2019), and further investigation of how to correct the Stokes drift to account for this could be highly relevant for practical applications. It may also be of interest to go beyond the confines of the third-order theory presented here, and to explore particle trajectories for solutions to the full Euler equations e.g. via conformal mapping techniques.
Appendix A On the non-uniqueness of Stokes expansion
The perturbation expansion associated with Stokes waves is often presented in textbooks as following a set sequence of steps. In fact, numerous choices are made, including whether to add linear terms in or to the potential, and how to treat constants appearing in the expansion. A useful discussion of some of these choices is found in (Svendsen, 2005, Ch. 8). We remark here on one aspect of this non-uniqueness that has received rather less attention (though see the comment in (Janssen, 2009, Appendix A)).
For a chosen linear wave amplitude there is a non-uniqueness in the Stokes’ expansion, which makes it necessary to specify the potential explicitly. For example, the following pairs of third-order solutions have the same linear amplitude
and
The former (47)–(48) is given by (Wehausen & Laitone, 1960, (27.25)), and appears naturally when the third order potential is set to zero. However, at each order solutions to the Laplace equation permit the addition of first-harmonic terms of the form . The latter (49)–(50) is the natural form obtained from the Zakharov formulation (see also (Janssen, 2009, (A.16)), (Gao et al., 2021, (6.7c))). Note that this flexibility in the perturbation expansion explains the seeming discrepancy identified by Zhang & Chen (1999) when matching these solutions.
Appendix B Coefficients of the higher-order potential
We provide here the coefficients of the second-order and third-order contributions to the potential from Section 2.3. The kernels appearing therein follow the usage of Krasitskii (1994).
Second-order coefficients are:
The third-order coefficients are:
As previously the shorthand notation is used to denote , and similarly for