### v2

# Wave-induced drift in third-order deep–water theory

Abstract
The goal of this work is to investigate particle motions beneath unidirectional, deep-water waves up to the third-order in nonlinearity. A particular focus is on the approximation known as Stokes drift, and how it relates to the particle kinematics as computed directly from the particle trajectory mapping. The reduced Hamiltonian formulation of Zakharov and Krasitskii serves as a convenient tool to separate the effects of weak nonlinearity, in particular the appearance of bound harmonics and the mutual corrections to the wave frequencies. By numerical integration of the particle trajectory mappings we are able to compute motions and resulting drift for sea-states with one, two and several harmonics. We find that the classical formulation provides a slight underestimate of the drift at the surface, and a slight overestimate at depth. Incorporating difference harmonic terms into the formulation yields an improved agreement with the drift obtained from nonlinear wave theories, particularly at greater depth. The consequences of this are explored for regular and irregular waves, as well as parametric wave spectra.
1 Introduction
Waves on the surface of water are a continuing source of fascination, as well as posing an important scientific and engineering challenge, with impacts on such diverse areas as offshore renewables and marine ecology. While we can observe the myriad forms of waves on the water surface, these shapes are only manifestations of what is going on below the surface: a transport of energy and, to a lesser extent, the movement of water particles themselves. Leonardo da Vinci already observed that the wave outruns the water, and early mathematical analyses of wave kinematics seemed to confirm this idea. Work in the early 19th century by Green showed that particle paths in linear deep water waves are approximately circular – in seeming agreement with the exact, deep-water solution derived by Gerstner – and the corresponding finite depth particle paths were shown to be approximately elliptical by Airy (Craik, 2004).
Nonlinear wave theories began to be explored by Russell, Kelland, and most prominently, Stokes, starting in earnest in the 1840s. It was the work of Stokes that first suggested that particle trajectories in periodic waves in deep water might not be closed, and gave rise to the idea of wave-induced motion now known as Stokes drift. In the intervening years the kinematics of nonlinear waves has received significant attention, and the concept of wave-induced drift has taken a prominent role in our understanding of transport in the ocean. Its consequences are still being explored in the movement of floating microplastics (van Sebille et al., 2020) and water-borne bacteria (Ge et al., 2012), and we are coming to develop a fuller picture of particle trajectories in nonlinear wave equations (Carter et al., 2020; Curtis et al., 2018; Ige & Kalisch, 2024).
Nevertheless, much remains unknown from both an experimental and theoretical perspective. Experiments are notoriously difficult, owing to the challenge of comparing a closed flume (with no net mass transport, and with the additional complication of a mechanical wavemaker) with the situation in the (open) ocean. Recent years have seen significant progress, from flume experiments on monochromatic wave trains (Umeyama, 2012; Paprota & Sulisz, 2018) which have been taken to higher and higher steepness (Grue & Kolaas, 2017) to several recent investigations of particle kinematics for wave groups and complex flows (Bjørnestad et al., 2021; van den Bremer et al., 2019).
Meanwhile, theoretical investigations must contend with the fact that the ubiquitous, Eulerian description of fluid motion – dispensing with particle identity in favour of a simpler treatment of velocity fields – necessarily obscures certain aspects of kinematics. While these difficulties can be circumvented for steady flows by the use of alternative formulations, including employing stream function and velocity potential as independent variables as originally suggested by Stokes, there is no easy fix for nonsteady, irregular wave fields. While the pre-eminence of Eulerian theory is being challenged by a resurgence of interest in Lagrangian wave mechanics, including higher-order approximate formulations for steady periodic waves (Clamond, 2007), standing waves (Chen & Hsu, 2009) and, recently, wave packets (Abrashkin & Pelinovsky, 2018; Pizzo et al., 2023), the vast majority of studies continue to employ the Eulerian perspective. The non-trivial task of comparing the two approaches (Fouques & Stansberg, 2008) should not be underestimated.
The aim of this work is to explore the theoretical particle trajectories – and drift – associated with nonlinear potential flow theories in infinite water depth. While there are recent mathematical results for the full, nonlinear problem showing that no closed particle paths exist in steady, periodic wave trains (Constantin, 2006; Henry, 2006; Okamoto & Shoji, 2012), these results are not quantitative and are difficult to generalise to nonsteady problems. We will take a ‘weakly nonlinear’ approach, and our starting point will be the Stokes drift formulation which arises in linear (1st order) water wave theory. While a rigorous proof that linear, steady, periodic water waves exhibit a forward drift in the direction of wave propagation (Constantin & Villari, 2008; Constantin et al., 2008) was recently found, the approximate results due to Stokes are still used in most applications. These were generalised to a spectrum of waves by Kenyon (1969), and the resulting formulation has been extensively explored for different spectral shapes (Webb & Fox-Kemper, 2011), for its sensitivity to the high-frequency spectral tail (Lenain & Pizzo, 2020), and new profiles have been developed for practical implementation (Breivik et al., 2016). Many more references can be found in the recent review by van den Bremer & Breivik (2018).
As the Stokes drift formulation comes about through a simplification of the linear water-wave theory, the differences between it and the linear theory are of interest. This also raises a natural question: what if we progress from linear theory to a second or third–order theory? We expect the higher-order theories to be more accurate, and to include important effects for steeper waves – such as the appearance of additional harmonics – which may have a bearing on comparisons with the Stokes drift. Indeed, these theories may suggest improvements to the Stokes drift derived from linear theory. However, with the exception of recent work of van den Bremer and Taylor (2016) focusing on wave groups it is difficult to find such comparisons in the literature.
For a successful comparison we need some way to compute the fluid velocity field , which drives the particle kinematics through the particle trajectory mapping We need a methodology that can deal with several Fourier modes, to simulate irregular seas, and which is accurate to a desired order of nonlinearity. For more than two Fourier modes the Lagrangian theory is currently limited to second order (Pierson, 1961), while important resonant effects in deep water occur only at third order. Consequently, and to increase the accessibility of the results, it seems expedient to work in Eulerian coordinates. Even here several options exist, including the use of traditional perturbation theory to obtain a solution to the water wave problem, which has been extended to third-order by Madsen & Fuhrman (2012). An alternative – which we employ – is to use the equivalent Hamiltonian expansion, which is due to Zakharov (1968) and for which the third-order solution has been explored by Zhang & Chen (1999) and Gao et al. (2021). The Hamiltonian expansion furnishes our velocity field and free surface to third-order, and these are the tools we use to examine
The remainder of this paper is structured as follows: in Section 2 we introduce the water wave problem and corresponding Hamiltonian. We give more details about the expansion that allows us to extract linear, second, and third-order terms in Section 2.1, discuss the neglect of amplitude evolution for our particle kinematics calculations in Section 2.2, and finally show how the free surface and potential are obtained to each order in Section 2.3. In Section 3 we consider monochromatic waves, which is the natural setting to introduce Stokes drift, and look at the consequences of higher-order theories on particle trajectories. Section 4 is devoted to bichromatic waves. These are the simplest type of non-steady wave-field, but because only two harmonics are involved the linear superposition is always periodic. Finally, Section 5 considers waves involving multiple lowest-order harmonics. A discussion of the work and some conclusions are presented in Section 6. Appendix A adds some comments on the non-uniqueness of the monochromatic Stokes wave solution at third-order, while Appendix B gives expressions for some coefficients appearing in Section 2.3. Appendix C explores the effects of slow amplitude evolution on particle paths and Stokes drift.
2 The water wave problem and Hamiltonian formulation
We begin with the inviscid, irrotational and unidirectional water wave problem in Eulerian variables. This can be written in terms of a fluid velocity potential and free surface as
The advantage of this formulation is that all nonlinearities have been moved to the boundary conditions, which therefore represent the principal challenge. We also note that, provided the water is deeper than about half a typical wavelength, the still-water depth can be taken to be infinite. This procedure, using
| (1d’) |
in place of (1d) simplifies calculations considerably. Despite the restriction to irrotational flow and the neglect of surface tension, compressibility and viscosity, these equations contain a huge range of interesting physics, including many types of wave phenomena. While the techniques we shall employ can be used without alteration for waves with a transverse () dependence, the additional freedom this entails means that we defer an exploration of directionally spread waves to future work.
During the late 1960s, efforts to find a Hamiltonian formulation of the water wave problem (1a)–(1d) came to fruition in work of Zakharov (1968) who found the Hamiltonian
which represents the total energy of the fluid. The domain of integration in can in practice depend on the specifics of the problem, and the lower boundary of the -integral can be chosen as if (1d’) is used in place of (1d).
The canonical variables for the Hamiltonian formulation are defined at the free surface. They are and and the canonical equations are thus
These equations are equivalent to the surface boundary conditions (1b), (1c), and details of the calculation can be understood most easily in Broer (1974). Additional background can also be found in the recent review by Stuhlmeier (2024).
It is worth noting at the outset that any study of kinematics is necessarily somewhat hampered by the Eulerian variables commonly used in wave research. The Eulerian focus on velocity fields as functions of fixed (laboratory) coordinates obscures the motions of fluid particles, which must be laboriously re-derived. Moreover, as we shall see below, an approximate Eulerian theory divorces the motions of the fluid and the free surface, with consequences for our ability to understand the full kinematics. The alternative choice of Lagrangian variables has many advantages (Pizzo et al., 2023; Blaser et al., 2024), but a general theory of nonlinear, irregular waves has yet to be developed, and the connections to the well-established Eulerian theory remain an topic of active exploration.
2.1 Asymptotic analysis and the water wave problem
The water wave problem, whether written as a system of partial differential equations and boundary conditions or using the Hamiltonian reformulation remains mathematically very complex. While the mathematical existence of periodic and solitary wave solutions has been established in ever more general settings (for example, the books by Constantin (2012) or Lannes (2013) offer a relatively recent overview of mathematical progress), much of what is known about qualitative and quantitative properties of ocean waves rests on an analysis of simplified equations.
These simplified equations are usually obtained through so-called asymptotic analysis or perturbation theory – a procedure of nondimensionalisation and scaling the equations, followed by the identification of small dimensionless parameters, and an asymptotic expansion of the variables in terms of the small parameters selected. In the case of water waves the most common small parameter is the wave steepness written either as the wave amplitude divided by the wavelength or multiplied by the wavenumber. A very thorough discussion of asymptotic analysis as applied to water waves can be found in the textbook by Johnson (1997).
Instead of expanding the water wave equations (1a)–(1d) in this manner (see Madsen & Fuhrman (2012) for a detailed, relevant account), it is also possible to expand the Hamiltonian directly. This procedure was carried out by Zakharov (1968) and, with more detail, Krasitskii (1994), and it has the advantage of enforcing structure that might otherwise be lost, such as that resulting equations conserve energy.
The first step in this procedure is to write the canonical equations in terms of the Fourier transformed variables and as
with denoting the complex conjugate. Then one defines a new pair of canonical variables and related to and via Fourier transforms as
where
is the frequency (in rad/s), and is the constant acceleration of gravity (taken to be 9.8 m/s2 in computations). As we will be interested in deep-water waves (, we note that the dispersion relation simplifies to
We will use this in all the computations that follow.
The subsequent steps in the expansion are somewhat lengthy, and are given by Krasitskii (1994). At this stage it is important to emphasise only one key point: at lowest order the problem is linear, so the superposition principle holds. This means that the Fourier amplitudes – which correspond essentially to the magnitudes of the terms introduced in (4)–(5) – do not change with time. If gravity is the sole restoring force, the surprising fact is that the Fourier amplitudes do not change with time at the next order, either. Only at third-order in nonlinearity, where resonance between quartets of gravity waves becomes possible, do we obtain an evolution equation. This equation gives the slow evolution of the complex amplitude
and is known as the Zakharov equation, with a kernel term given in (Krasitskii, 1994, (3.5)). We have introduced subscript notation to denote wavenumbers, so that, for example, the delta-function is a shorthand for In principle can be either a scalar or a vector wavenumber, although our computations will deal only with the scalar case for simplicity.
If is known then the canonical variable can be recovered via
where we have written terms up to third order only. This expression is given in (Krasitskii, 1994, (2.17)), as are the relevant interaction kernels . Finally, once is known it is possible to recover and up to the desired order by means of (4)–(5).
2.2 Constant magnitude approximation
The procedure outlined above would seem to suggest that we first need to solve the complicated integro-differential Zakharov equation (8). In fact, for certain situations of interest we can circumvent this step. The key insight is that the evolution described by the Zakharov equation – being associated with cubic nonlinearities – occurs on the slow timescale , where is a typical wave steepness and is a typical wave period. For processes taking place on faster time-scales, it may be appropriate to consider the complex amplitudes to be constant.
This is indeed the approach we will adopt. There is, however, one important addition: for certain wavenumber combinations the Zakharov equation (8) can be solved explicitly. In particular, for all symmetric quartets which satisfy or the equation reduces to a system of one or two ordinary differential equations with constant-amplitude solutions, as shown in (Mei et al., 2018, Ch. 14). Such solutions represent nonlinear corrections to the frequency of the waves, and were first found by Stokes for a single wave mode. For arbitrarily many modes, or a continuous spectrum, it is possible to write these corrections in a compact form (Stuhlmeier & Stiassnie, 2019).
Incorporating these corrections means that the complex amplitudes have constant magnitudes but time-dependent phases
The terms in the argument of the exponential reduce to , the well-known Stokes’ correction, when only a single wave-mode is present, or the mutual frequency correction found by Longuet-Higgins & Phillips (1962) for two modes. The form (10) now contains all the mutual dispersion corrections entailed by the cubically nonlinear water wave theory. It has been demonstrated that taking these corrections into account allows one to predict the short-time propagation of laboratory waves with high accuracy without the need to solve for the evolution of the slow-time variables (Stuhlmeier & Stiassnie, 2021; Galvagno et al., 2021; Meisner et al., 2023), showing that the constant magnitude approximation is well-justified for short times. Of course, under specific circumstances – for example with a sufficiently narrow spectrum, or close to modulation instability – we may have significant energy transfer (Andrade & Stuhlmeier, 2023). However, given our focus on particle paths (with a characteristic time on the order of one period) the amplitude evolution is negligible (see also the discussion in Stuhlmeier & Stiassnie (2019) and Appendix A therein).
If we sum up the consequences of weak-nonlinearity for surface gravity waves in deep water as
1.
dispersion corrections,
2.
changes to wave shape (bound harmonics),
3.
energy exchange between modes,
we will account for 1–2 while neglecting the slow effects 3. Appendix C provides some further discussion of the neglect of amplitude evolution, along with several examples.
2.3 Recovery of the third-order solution
The Hamiltonian formulation makes clear how the free surface and potential can be recovered from the canonical variables and up to a given order. In practice, discrete wavenumbers are considered, which means that the integrals in (9) can be simplified by the substitution
Here is unknown in principle, and in practice comes from a constant magnitude approximation to (8) as discussed in Section 2.2.
Because the free surface (4) is recovered directly from substitution of (9), it is clear that we can distinguish linear, quadratic and cubic terms in based on the respective terms in (9). The same is not true of , as we shall see below. Taking the inverse Fourier transform yields
Here we have substituted with amplitude and phase and employ the compact notation Clearly this yields the expected sinusoidal free surface elevation upon defining the amplitude
The nonlinear corrected frequency is given by
and for unidirectional waves in infinite depth – which will be our focus – we can simplify the kernels to
The quadratic components of the free surface elevation can be obtained in analogous manner, with slightly more in the way of algebra required to resolve all delta-functions. We find
The shorthand notation is used to denote
We can write the cubic components of the free surface elevation as
The velocity potential at the free surface, , is obtained from (5) via an identical procedure.
The corrections contained in the second-order and third-order free surface elevation essentially change the wave shape. The resulting harmonics of the form behave as waves with wavenumber and frequency but because these sums and differences do not satisfy the dispersion relation they are referred to as “bound waves”, in contrast to the free waves of (12). The well-known impact of these bound waves on sharpening the crests and flattening the troughs is shown in Figure 1 for a bichromatic sea, together with the effect of the third-order frequency correction (13) on the wave phases. We also note the important effect of the difference harmonic terms in creating a “set-down” below the crest, see van den Bremer et al. (2019).
2.3.1 Recovery of the bulk potential
The reformulation of the problem in surface variables is very convenient from many perspectives, but is something of a disadvantage when we wish to recover information about the bulk fluid below the free surface. To obtain this we start with the general solution of the Laplace equation which decays for infinite depth
the exact analogue of (Krasitskii, 1994, (4.1)).
We relate the bulk potential and its surface trace to one another in terms of an expansion. Taking the problem in infinite depth, we need to solve
which we do by assuming the free surface elevation is small (), and then inserting this expansion into the boundary condition (18), so that
This gives rise to a hierarchy of problems in physical space, whose successive solution yields
with
With and known from the formulation in (4)–(5) we can now recover the bulk potential at each order desired. This procedure is also detailed in (Janssen, 2004, below (4.7)), (Krasitskii, 1994, p.15) and (Zakharov, 1968, (1.8)), although care should be taken in the symmetrisation of (21).
The linear contribution consists simply of the first term in (20)
The quadratic contribution is
where we need to employ the linear and quadratic parts of and (and products thereof) in the formulation. The coefficients are given in Appendix B. Up to second order, for unidirectional waves in deep water the potential reduces to the simple expression
This follows also from Dalzell (1999).
The cubic contribution to the bulk potential is
The coefficients are given in Appendix B. Unfortunately efforts to find a compact simplification the third order have failed except in the case of a single monochromatic wave, where the analytical expression
It should be emphasised that the bulk potential thus obtained is essentially equivalent (except for the non-uniqueness at a given order mentioned in Appendix A) to the potential found through direct perturbation expansion of the governing equations (1a)–(1d’). In particular, as we shall see below, this means that the fluid domain at each order is the half-space owing to the transfer of the boundary conditions.
3 Monochromatic waves
The mathematically simplest type of wave motion, and the one about which we know the most, is that of steady, periodic waves known as Stokes waves. For such waves, symmetric about a crest and propagating without change of form, it can be proven that no closed particle trajectories exist (Constantin, 2006), and explicit calculations of the surface drift have been given by Longuet-Higgins (1987), up to and including the wave of greatest height.
Because the wave-form is steady and the fluid motion is confined to the -plane, a variety of powerful mathematical techniques exist for tackling the problem, including the use of the velocity potential and stream function as independent variables (Stokes, 2009). Moreover, the convergence of perturbative solutions for such waves has been established since the work of Levi-Civita (1925), and in practice such solutions have been computed to extremely high order (Schwartz, 1974). This wealth of prior results means that we do not strictly need the Hamiltonian expansion developed in Section 2; however we shall attempt to situate the 3rd order theory and its predictions among other results for monochromatic waves.
Thus, if we fix the wave length (or wavenumber) and the wave height the successive inclusion of higher order terms brings us ever closer to the exact solution. This appears to be true also when we employ the equivalent Lagrangian formulation of fluid mechanics (Clamond, 2007), albeit without the accompanying analytical theory proving convergence. Some care must be taken in the selection of expansion parameter (in which the “order” is measured), as discussed by Cokelet (1977) and Fenton (1985), about which more will be said below.
Although these steady, periodic waves are theoretically exceptional and experimentally difficult to realise, the wealth of accumulated understanding makes them an ideal starting point, and allows for comparisons that will stand us in good stead for later, unsteady problems. The textbook monochromatic solution (Dean & Dalrymple, 1991) to the linearised version of equations (1a)–(1d’) is
where and
is the linear dispersion relation. Here and
The particle trajectory mapping is the system of ordinary differential equations
It is worth noting that this system has only a tenuous connection to the linear (or, later, weakly nonlinear) theory, but rather attempts to restore the link between Lagrangian and Eulerian descriptions of the fluid motion. In addition, it is an unwelcome surprise that inserting the linear potential yields a nonlinear system of differential equations
While this system cannot be solved analytically it has been shown rigorously that its particle trajectories are not closed (Constantin & Villari, 2008). The result is qualitative, and an analytical treatment that yields the drift quantitatively even in this simple case appears out of reach.
In lieu of this, two alternatives remain: approximation or numerical solution. The former approach is found in most textbooks on water waves, and begins with a Taylor expansion of the fluid velocity field about an initial position . Keeping only the lowest order terms in the Taylor expansion yields, upon integration, the circular particle trajectories first found by Green in 1839 (Craik, 2004). Inserting that solution into the second-order Taylor expansion yields another explicit system of ODEs
Integration of this system yields the familiar result that particles move forward with a Stokes drift velocity
This Stokes drift is nominally a second order – and therefore nonlinear – quantity in the small wave amplitude , but is derived from linear theory (25). This important reminder can also be found in many textbooks, such as Dean & Dalrymple (1991).
Figure 2 compares the solution of the linear particle trajectory mapping (28)–(29) with that of its approximation (30)–(31). Following Longuet-Higgins (1987), we define the Lagrangian period as
where is the Eulerian period, is the phase speed of the wave, and is the Stokes drift at a given depth. In the approximate model, shown with dashed curves in Figure 2, after one Eulerian period each particle has moved a uniform amount to the right (diamond markers). By contrast, after one Lagrangian period, calculated using (32), the linear theory yields a forward drift that depends on the initial position (square markers), and which is sometimes smaller and sometimes larger than the Stokes drift. This dependence of the drift on the phase was also observed experimentally by Grue & Kolaas (2020).
More importantly, Figure 2 emphasises a crucial blind-spot of the linear theory: the fluid velocity field is defined only below Thus even a particle starting initially at trough level – such as those on the right side of Figure 2 – will enter a region of the -plane where the velocities are undefined (indeed, they may also exceed the crest height, as can be seen in the red trajectory on the far right; we should not make too much of this fact, since the free surface in linear (Eulerian) theory is not a streamline nor is it composed of fluid particles).
An evaluation of the velocity field above amounts to an extrapolation of the linear theory, a procedure which is known to overestimate velocities near the surface (Johannessen, 2010). One possibility to obtain surface drift values is integration of the horizontal surface velocity (see Grue & Kolaas (2020), who use this to 2nd order to compute the Lagrangian period) where the right-hand side can be approximated as
As mentioned, we do not strictly require the Hamiltonian expansion for monochromatic waves, and in practice we make use of asymptotic expansions up to fifth order in which were derived by Stokes for deep water and are reviewed in the comprehensive recent paper by Zhao & Liu (2022).
The results are shown in Figure 3, which compares the formula (32) (label SD) with weakly nonlinear approximations up to 5th order (omitting the 1st order, which is effectively unusable), as well as the exact surface drift obtained by Longuet-Higgins (1987) (label LH). Taking 1/m, the horizontal axis denotes half the actual crest-to-trough height, and not the leading-order amplitude which is used in the perturbation expansion, and which must be adjusted every odd order to obtain a desired value of (see discussion in Appendix A). Without such an adjustment the wave height grows with the order of the expansion, as does the surface drift.
Figure 3 shows that for waves of low steepness (below ) the differences in the formulations are essentially negligible. The Stokes drift formula (32) gives a drift that is slightly too small (as shown by Longuet-Higgins (1987) and complementary work using the Lagrangian formulation, such as Clamond (2007)), but higher-order theories provide good agreement up to very high steepness (though the trend clearly does not continue to the steepest wave with ). We note that dispersion corrections appearing at the 3rd and 5th order have the effect of decreasing surface drift slightly, while the seemingly excellent agreement between the 2nd order approximation and the exact solution at slopes above should be considered essentially accidental.
The depth-dependent Stokes drift can be obtained from integrating the particle trajectory mapping, provided particles do not cross the still-water level – otherwise, as mentioned above, extrapolating the terms can lead to spuriously high velocities or even divergence of the trajectories. Because the forward drift of a particle is phase-dependent, this means averaging over the phases of particles lying at an initial level In Figure 4, 50 points covering the wave phase are used at each depth to compute the effective Stokes drift. We compare results obtained from 1st, 3rd and 5th order velocity fields, as well as a 4th order Stokes drift from the Lagrangian obtained by Blaser et al. (2025) (see also Clamond (2007)), and which we shall return to for unsteady cases below.
It is noteworthy that the integration of the 1st order velocity field (red curves, label 1) gives a significant overestimate of the Lagrangian drift with depth, while the approximation (32) based on that same velocity field (blue curves, label SD) gives much better agreement with both the 4th order Lagrangian result and higher-order Eulerian velocity fields. Only for high steepness and shown in panels (b)–(c) is the approximate Stokes drift noticeably different from the higher-order solutions, although its asymptotics at large depth hew close to the 1st order theory, while the higher-order solutions have a somewhat different behaviour at large depth (see inset in panel (c)).
In the formulation of the velocity potential we are using here (corresponding to (44a)) the only difference in velocity potential between the 1st and 3rd order is only the inclusion of nonlinear dispersion, so that waves travel with a characteristic coordinate for
No additional harmonics appear until 4th order, so the difference between curves 1 and 3 in Figure 4 is entirely a consequence of this dispersion correction. Indeed, the procedure used to derive (32) can be applied without alteration to the 3rd order potential, yielding
Because it is clear that the drift velocity (35) is generally smaller than (32), as observed also from the particle trajectories.
We can summarise the results for monochromatic waves as follows: the approximate Stokes drift formula (32) gives results that are somewhat too small at the surface, and somewhat too large at depth. The differences for waves of moderate steepness, however, are not sizeable, as borne out also by experimental work (Paprota & Sulisz, 2018). In our description, the only effects of nonlinearity are in adjusting the frequency (which has consequences for the entire flow) and in the appearance of higher-harmonics (whose effect is confined close to the surface due to the -terms). We shall now see how these effects appear in unsteady (bichromatic) wave trains.
4 Bichromatic waves
The simplest “irregular” waves, corresponding to unsteady flows, are those that consist of two harmonics and at first order, and are therefore sometimes termed “bichromatic” waves. The second order solution then contains sum and difference terms of any two of these, while the third order contains sum and difference terms of any three, with an attendant increase in complexity. Up to second order the solution is contained in the work of Dalzell (1999), while the third order solution for bichromatic waves appears first in the work of Madsen & Fuhman (2006). We make use of the equivalent formulation of Section 2.3.
Up to second order the velocity potential can be written as the remarkably simple expression
where without loss of generality, and the sum harmonic vanishes for unidirectional deep water waves (the same is not true of the free surface elevation, whose prominent sum harmonics can be seen in Figure 1).
We show the free-surface time series of one such bichromatic solution in Figure 5(a), where the inclusion of 2nd and 3rd order terms can be seen to have only a minor effect. This is not the case when considering the particle paths, which exhibit notable differences between linear and higher-order theories, particularly at depth. Figure 5(b) shows the paths of particles close to the surface with initial position (marked by ) and different final positions (marked with , and in the respective colours).
Note that Figures 5(b),(d) are simply obtained by integrating where denotes the velocity at a given order ( or 3). The relative importance of each term can be assessed by integrating separately for each , which yields Figures 5(c),(e). From these it is clear that the principal contribution at the surface comes from the 1st order velocities, with the principal contribution at depth coming from 2nd order velocities. The 3rd order contributions (in yellow) are barely visible near in panel (e).
The second order solution (36) lends itself to a calculation of a modified Stokes drift in a rather straightforward manner: calculate from (36), Taylor expand about a point , keep only the constant terms in the Taylor expansion and calculate approximate trajectories , subsequently substitute these trajectories into the linear terms in the Taylor expansion and integrate the resulting particle trajectory equations in time. The result of this procedure is
which consists of the sum of the Stokes drifts of each mode separately (I) as well as a new term222While this paper was under review, an article by Liao & Zou (2025) appeared which also derives this term and the extension (41), see (C11) in their Appendix C. (denoted (II)) that arises from the difference harmonics (recall that we assume ). Just as the Stokes drift derived from linear theory includes quadratic terms, so the modified Stokes drift derived from second order theory now includes (formally) quartic terms.
While overall particle motion decreases as we descend into the fluid, there are regions where the 2nd order contribution to Stokes’ drift has a sizable effect (see Figure 6(b)) and in fact dominates the 1st order contribution (I), as shown around in Figure 6(c).
Figure 6 shows the terms (I) and (II) of (37) at three depths: (a) , (b) and (c) for a fixed wave 1/m. It is simplest to analyse the relative importance of the terms by setting and varying We note that the contribution (II) at the top of the fluid domain grows monotonically with , as can be observed in Figure 6(a) (note that the steepness of both modes is fixed at ). We also note that there is a minimum of the total 2nd order Stokes drift at the surface. By contrast, for the quartic term (II) has a single maximum and subsequently decays, as shown in Figure 6(b)–(c). As we descend into the fluid, this maximum moves towards smaller values of
The principal difficulty we encounter – for the first time with bichromatic waves – is connected to the unsteadiness of the flow field and the lack of a unique solution towards which different expansions are known to converge. Unlike the monochromatic waves considered in Section 3, we cannot simply compare waves of the same length and height, which differ geometrically only in the curvature of the profile between crest and trough. The problem is compounded at third order by the appearance of an asymmetric frequency correction (for ) first found by Longuet-Higgins & Phillips (1962)
This means that bichromatic waves at third-order (and 5th, 7th, and so on) will be out of phase with their lower order counterparts, as seen in Figure 1.
Compounded by a paucity of results (theoretical or experimental) on drift associated with bichromatic waves, this makes comparisons difficult. However, we can employ the methodology tested for monochromatic waves in Section 4 to obtain results for bichromatic waves at the surface and at depth, and compare these to the approximate Stokes drift derived in (37).
To obtain drift at the surface we integrate the equation for the horizontal particle position where the velocity field is given by Taylor expansion about to a given order. The 2nd order horizontal velocity can immediately be read off from (36), while the 3rd order velocity, for which a simple algebraic expression has not been found, is obtained from the Hamiltonian expansion of Section 2.
Figure 7 compares the 2nd and 3rd order theory with the classical Stokes drift (SD I) and the modified Stokes drift (SD I+II). In panel (a), where the two waves are close in frequency, we expect the contributions of the terms (II) to be negligible, and indeed the two formulations largely coincide. For more widely separated waves and rad/s, shown in panel (b) the difference harmonic contributions of (II) are more prominent. In both cases the 2nd and 3rd order theories give an enhanced drift, which is consistent with the results obtained for monochromatic waves and shown in Figure 3.
We have also compared with the 4th order drift obtained from the 3rd order bichromatic solution in Lagrangian variables obtained in Appendix B of Blaser et al. (2025), which generalises Pierson’s (1961) classical 2nd order result. Unfortunately, bichromatic waves in Eulerian and Lagrangian coordinates are not the same; rather, the wave profiles are geometrically distinct, with a phase-shift that occurs between successive crests of both solutions, as already noted by Fouques & Stansberg (2008)333This phase shift is not due to differences in dispersion relation, as might be hypothesised from Pierson (1961), since Blaser et al (2025) recover the Eulerian dispersion correction (38)–(39) at 3rd order.. Whether it is possible to adapt the Lagrangian perturbation expansion to match the Eulerian solution (or vice versa) remains an open question. For well-separated frequencies (panel (b)) the Eulerian and Lagrangian solution remain remarkably close, while for two nearby frequencies (panel (a)) the notable difference grows with wave slope – and is visible also when comparing the free surface plots of the two solutions.
Below the surface, exactly as shown for monochromatic waves, it is possible to use the particle trajectories obtained by integrating the Eulerian velocity field to obtain the Stokes drift. This is put to the test in Figure 8, where we select a bichromatic wave train with significant subharmonic drift as per (37), namely and 1/m. This corresponds to the circles shown in Figure 6.
Figure 8 compares 1st, 2nd and 3rd order theory with the formulations (I) and (I)+(II) of (37), as well as the 4th order Lagrangian drift of Blaser et al. (2025). In fact, the particle paths obtained from integrating 1st order theory coincide completely with the classical Stokes drift (SD I). On the other hand, there is very good agreement between 2nd and 3rd order theories, the improved Stokes drift formulation (SD I+II) and the fourth order Lagrangian theory (L4). The same comparison for a pair of more widely separated wavenumbers (such as and 1/m, not shown) would demonstrate that all six curves essentially coincide, as might be predicted from Figure 6 (noting that there are no notable flows induced by the difference harmonic).
4.1 Difference harmonics and particle drift
In general, we see the importance of leading order effects near the surface of bichromatic waves, while second order effects become more prominent at depth. This is no surprise, and follows from the structure of the difference harmonic terms themselves. As a prelude to our work on wave groups it is expedient to investigate their role in bichromatic seas.
The first consideration should be of the velocity field which drives the particle trajectory mapping. We consider the waves studied in Figure 6 with and 1/m, and calculate horizontal velocities directly from (36). Indeed, Figure 5 makes clear that the 3rd order velocity field – though difficult to calculate – makes an essentially negligible contribution. In fact, the only material contributions of the 3rd order theory are due to the dispersion correction (38)–(39), as can be seen by comparing panels (b) and (d) with panels (c) and (e) in Figure 5. These dispersion corrections do not affect the instantaneous velocity field, only its time evolution.
When only the first order contributions are retained we find the familiar pattern of forward velocity below the crests and backward velocity below the troughs, with an exponential fall-off in depth as depicted in the top panel of Figure 9. If instead we focus only on the second-order contributions , the situation looks quite different. The exponential decay in velocities is slower because is larger than either or for and we find a negative horizontal velocity under the centre of the bichromatic wave train, as shown in the bottom panel of Figure 9. The full second-order picture is a sum of the first and second order terms, shown in the middle panel of the same figure, and it is clear that the first order dominates near the surface while the second-order difference harmonic dominates at depth. This can be observed clearly by considering the zero contour lines of the horizontal velocity, shown in black for each case.
The generic situation for particle paths in a bichromatic wave train is thus as follows: close to the surface the first order contributions dominate, giving the curlicue particle paths observed in Figure 5(b). The effect of difference harmonics is to retard slightly the forward motion of the particles, due to the slow moving (in deep water the group velocity is half the phase velocity) region of backward (negative) velocity under the highest crests observed in Figure 5(e).
Deeper into the fluid column the effect of the first order free-modes becomes negligible. The fluid motion is dominated by the difference harmonic term , and the particle paths behave essentially as those of a monochromatic wave with that wavenumber. This is clearly illustrated in Figure 5(e), where we see that the modes and which make up the first-order theory are only a small perturbation on the displacement due to the difference harmonic term.
It is worth pointing out the existence of a region beneath the bichromatic wave train where the particle displacement is largely against the direction of wave propagation, as seen between in Figure 5(e) (see markers ’+’). This should not be surprising, since it simply reflects the effective motion of the difference harmonic term as shown in panel (d). Despite the existence of this ‘return flow’ – which we shall encounter again in our discussion of wave groups – the net particle motion remains forward at every depth in the fluid, as shown in Figure 6 and Figure 8 below. Because the solution is periodic, over many periods the net effect of difference harmonics is to enhance forward drift, as shown in Figure 8.
5 Multichromatic waves
When more than two Fourier modes are necessary to describe the linear problem, we call such a configuration multichromatic. The second order contains sums and differences of any two modes, while the third order contains sums and differences of any three. This means that significant algebra can be involved in calculating higher-order solutions.
5.1 Focused linear wave groups
In treating waves with an increasing number of linear harmonic constituents we also have a plethora of parameters that can be tuned. One case of interest, from both an experimental and a theoretical standpoint is that of a (focused) wave group. Linear focusing is quite straightforward: if our free surface is a superposition of sinusoids as per (25), then writing
readily gives a wave train that focuses – i.e. all components come into phase – at . Pulse-like wave groups can be created by suitable tuning of the amplitude spectrum
One possible choice is a Gaussian amplitude spectrum of the form
An example of a wave train with such an amplitude spectrum is shown in Figure 10(a), where we have chosen equidistant modes between and 1/m, , and The focusing location is chosen to be , and the focus occurs at a crest. Following Blaser et al. (2025) we define
as the maximum slope for the linear focused wave, and use this as a measure of nonlinearity going forward.
The particle behaviour in Figure 10(b),(d) is in many ways similar to simpler, bichromatic wave group shown in Figure 5. Towards the top of the fluid column, 1st order theory dominates, while at depth the effect of the in-phase subharmonics associated with higher-order theories clearly accounts for the bulk of the particle motion. In order to further illustrate this point, panels (c) and (e) show only the horizontal component of the particle position as a function of time , as the wave group passes over the initial position. Towards the surface (c) there is nearly no motion until s, while at depth we observe a steady forward drift in the positive -direction. Under the centre of the group, as expected, we see a strong forward drift towards the surface (panel c) and backward (return) flow at depth (panel e).
This effect can be seen even more clearly in Figure 11, which illustrates (for a single particle) the depth dependence of the horizontal motion. The (leading order) free surface of a focused Gaussian wave train with 1/m is shown in panel (a), in both the crest focused (blue, solid curve) and trough focused (red, dash-dotted curve) case. Panel (b) shows the horizontal displacement of a particle as the wave group passes it.
Interestingly, there is a marked difference in the return flow between crest and trough focused cases, with even 1st order theory showing some return flow for trough focused waves (dash-dotted curves). Such (localised) return flows are not captured by Stokes drift approximations, whether based simply on the linear theory or incorporating second order effects via the extension of (37)
Indeed, the effects shown in Figure 11 are highly phase-dependent, and rely on suitable choice of position in the spatio-temporal evolution of the group, while the approximate formulae (41) can only capture an averaged drift.
As we have done for monochromatic and bichromatic waves, we can apply 2nd and 3rd order, deep-water theory directly to calculate the surface drift of focused wave groups. The results of such a computation are reported in Figure 12, which is the counterpart of Figures 3 and 7. In this case, we have initialised our wave group with a central frequency and using modes have defined the amplitudes where the frequencies are chosen via for bandwidth and are the wavenumbers obtained from the linear dispersion relation. This is precisely the formulation used by Blaser et al. (2025), and allows us to compare our results with their theoretical and numerical values.
Following the trends for monochromatic and bichromatic waves, we note that the lowest order Stokes drift formulation (SD I, which treats the Stokes drift as a sum of the drifts of each individual mode (32)) gives the smallest value. The 2nd and 3rd order formulations show good agreement with both laboratory experiments () and fully nonlinear simulations () up to rather high steepness , even though our formulation neglects amplitude evolution which is included in the numerical simulations (see also the discussion in Appendix C).
5.2 Random multichromatic waves
Random waves are the closest mathematical idealisation to the waves we might typically find in the ocean: the phases of one Fourier component are not correlated with the next, and the resulting pattern – while occasionally exhibiting a distinct group – is highly irregular. Random phases at lowest order also means that sum and difference phases at higher orders are random, due to a neglect of slow phase-coherence effects associated with nonlinear coupling (Andrade & Stuhlmeier, 2023).
Where focused wave groups exhibit localised features, such as a clear division between flows in the direction of propagation at the surface and return flows at depth, – which have been well described by using wave-envelope formulations by van den Bremer and coworkers (Van Den Bremer & Taylor, 2016; van den Bremer & Breivik, 2018; van den Bremer et al., 2019; Calvert et al., 2019) – the lack of clear “groupiness” in random seas means that we should expect these localised effects to be averaged out. The advantage is that we can test whether (on average) the Stokes drift formulations developed in (41) are suitable, since they clearly cannot capture spatio-temporally localised flows around a group itself (all terms are positive).
Figure 13 shows the results of computing the wave-induced particle displacement for 200 realisations of a random sea surface. The amplitude spectrum (40) is retained, and the same parameters 1/m and as in Figure 11 are used, but rather than assigning the phases so as to achieve linear focusing at a specific location, these are randomly generated for each realisation. Blue curves denote first order theory, red curves second order, and yellow curves third order, with the prominent, darker curves representing the average over all realisations for a given order. In each case the integration is for approximately 80 peak periods, which is long enough to clearly distinguish the particle drift.
With a peak wavenumber 1/m, most of the motion due to the free-wave constituents is filtered out at a depth of m. The linear theory gives a very narrow range of displacements, with the realisations in red barely showing deviation from the average, and agreeing well with the linear Stokes drift (SD I). Second and third order displacements are quite large in some realisations, depending on the phase relationships involved. However, as expected, there is no systematic return flow at any depth for random waves, because such (spatially and temporally) localised flows are averaged out to yield the net forward drift which we observe in Figure 13. Moreover, all theories approach zero from above with depth, exhibiting the same ordering we have hitherto observed: linear theory predicting the smallest drift followed by second and third order. The correction (II) in (41) is found to give a useful improvement over using only the classical Stokes drift term (I).
5.3 Parametric wave spectra
The comparisons of the previous sections, based on explicit computation at different orders, support the inclusion of quartic difference harmonic terms in calculations of the Stokes drift throughout the water column. For many applications in ocean science and engineering it is advantageous to present such a result in terms of energy rather than amplitude spectra, and to explore its consequences for some common parametric spectral shapes.
For a continuous wavenumber spectrum , we make the identification with the modal amplitudes where the uniform grid spacing is assumed to tend to zero. With this assumption, is is straightforward to rewrite the discrete multi-chromatic Stokes drift with second order contributions as
where the first term is identical to that found by Kenyon (1969) and the second term comes from the inclusion of second-order terms in deep water.
Numerous critical discussions have focused on the key role of the high-frequency cut-off on the Stokes drift near the surface, including Breivik et al. (2014) and more recent work by Clarke & Van Gorder (2018) and Lenain & Pizzo (2020). As Clarke and van Gorder (2018) point out, if the Stokes drift is rewritten in terms of frequency and the high-frequency tail of the spectrum is such that , then we have a divergent expression for . Their proposed resolution, introducing a wave breaking frequency as an upper limit for spectral Stokes drift calculations was found to be reasonable for the subsurface Stokes drift by Lenain & Pizzo (2020), but we will employ the maximum wavenumber suggested by Lenain & Melville (2017) and used by Lenain & Pizzo (2020).
As a first comparison, we consider the Phillips spectrum for the equilibrium range of wind-generated waves, given in terms of radian frequency as
where we take as in Breivik et al. (2014). We relate the peak frequency to a friction velocity (see Holthuijsen (2007)) in order to establish as above, and show the results in Figure 14.
Panel (a) shows the Stokes drift near the surface using either the first term in (42) only (SD I) or both terms in the same equation (SD I+II). The difference between the two is sizable, but strongly dependent on the choice of high-frequency cut-off This tracks with the very pertinent discussion in Lenain & Pizzo (2020) (therein they report an underestimation of the Stokes drift of up to 50% from failure to account for high frequencies, but they do not consider the effect of difference harmonics). This difference persists as we descend into the fluid, and a noticeable difference on the order of 1 % is visible at depths below one peak wavelength in panel (b). Defining the Stokes transport as
we find nearly a 9% change in total Stokes transport. If we calculate the Stokes transport only below of the -folding depth of the peak frequency, we still observe a change of approximately 2% owing to the inclusion of difference harmonic terms (II).
The Phillips spectrum describes only the high-frequency tail, so it is of interest to consider a more realistic parametric spectrum: we next consider a Pierson-Moskowitz (PM) spectrum, corresponding to a fully developed under a constant wind over unlimited fetch. The case shown in Figure 15 corresponds to a wind speeds of 10 m/s (with corresponding ), and peak frequency of 0.17 Hz. Blue curves show the formula of Kenyon (1969) (corresponding to the first term in (42), labelled SD I) and red curves show both terms of (42) (labelled SD I+II). For reference, we note that the high-frequency cut-off obtained from is 6.3 We note also that it is possible to cut off the low frequencies, and evaluate the outer integrals in (42) starting from instead of 0 without materially affecting the results, as suggested by Breivik et al. (2016); Lenain & Pizzo (2020).
The differences in the formulations are clearly visible, both near the surface and at depth. At the surface, we find increases of surface drift in excess of 40% when difference harmonic terms are accounted for, reducing to just less than 2% at half the -folding depth of the peak wavenumber Interestingly, starker differences reappear at greater depths, as shown in panel (b) which highlights the different asymptotics of the two formulations at depths of more than one peak wavelength. We find that keeping both terms in (42) leads to an increase in Stokes transport of nearly 3.3 % for m/s. The majority of this difference comes from the near-surface transport, with transport below the peak -folding depth accounting for a difference of only 0.1%. The same analysis can be carried out for JONSWAP spectral shapes, with qualitatively very similar results to those found in Figure 15.
We can also compare the Stokes drift formulations of equation (42) for other spectral shapes, such as swell-wave spectra described by Ochi & Hubble (1976). In Figure 16 we show the most probable Ochi-Hubble shape corresponding to a significant wave-height m, and the two Stokes drift calculations. The surface drift shows a change in excess of 20 % when subharmonic terms (II) are included, and the Stokes transport is slightly more than 1% larger over the entire water column.
6 Discussion and conclusions
The wave induced drift of particles has attracted increasing attention in recent years, particularly in connection with the transport of maritime pollution. Nevertheless, explorations of the underlying hydrodynamics are not widespread. Notable exceptions are the recent articles by Van Den Bremer & Taylor (2016) and the follow-up study by van den Bremer et al. (2019), which focused on the transport associated with wave groups to second order. Our aim has been to explore particle drift systematically using the first, second, and third order theory, working our way from monochromatic waves to random seas.
Monochromatic waves – or progressive Stokes waves – are extremely well studied, and therefore serve as a useful test case for any development of nonlinear wave theory. They allow us to compare with exact solutions at the surface (Longuet-Higgins, 1987) as well as high-order approximate solutions in Eulerian (Zhao & Liu, 2022) and Lagrangian (Clamond, 2007) variables. The ease of defining a monochromatic wave – in the absence of an Eulerian current, which is the case for all the waves we consider, only an amplitude and a period are needed (Fenton, 1990) – means that approaches at each order can be directly compared to one another. Such comparisons show that the classical Stokes drift formula is remarkably accurate, at least when compared to other inviscid, irrotational theoretical results. Some differences do emerge for waves of large steepness, where the Stokes drift formula derived from linear wave theory tends to underestimate drift at the surface and overestimate it at depth.
With a view towards understanding more complex patterns of waves (and wave groups) on the open sea, we have subsequently studied the unsteady flow associated with bichromatic waves, generated at lowest order by two Fourier harmonics. Unfortunately, very few experiments have been carried out on bichromatic waves, and available results such as those by Westhuis et al. (2001) focus on the free surface elevation and the development of instabilities rather than induced flows. Other comparisons of horizontal velocities in bichromatic waves (Stansberg et al., 2008) rely on numerical simulations only. The problem is compounded by the fact that bichromatic waves are not themselves well-defined, since the inclusion of higher-order effects leads to a phase shift between the modes. This makes comparisons between different theories fraught with difficulty, and despite the fact that the bichromatic wave train is an exact solution to the 3rd order water wave problem Mei et al. (2018) no exact numerical solutions and attendant kinematics appear to be available in the literature.
That said, our investigations are able to shed some light on the flows induced by bichromatic waves. It is well known that the water column acts like a filter, first suppressing the highest frequencies and filtering out ever lower frequencies with increasing depth. Therefore the particle motion of monochromatic waves is typically considered negligible half a wavelength below the surface. It is this filtering effect on linear, free-wave constituents that leads to the dramatic decrease of drift in the linear theory for bichromatic waves. However, nonlinearity introduces much longer bound-waves which become dominant at depth, as was recognised in the landmark work of Longuet-Higgins & Stewart (1964). Beneath the centre of a wave group we find that these can induce sizeable localised return flows. However, their overall contribution is to increase the drift in the direction of wave propagation versus linear theory. Third-order contributions lead to additional bound waves, but also to dispersion corrections, which tend to retard the drift slightly.
In the bichromatic case the potential at second order includes new – difference – harmonics, in contrast to the monochromatic potential, which is identical at first and second order. This provides an impetus to update the Stokes drift formula, simply by using the second order potential in the Taylor series expansion in place of the first order. The new term – which captures the Stokes drift of second-order difference harmonics, can make a substantial contribution to the depth dependent drift depending on the wavenumbers involved. This contribution is also captured when simply integrating the particle trajectory ODEs in second- or third-order theory, as well as when using a fourth order Lagrangian drift due to Blaser et al. (2025). This corroborates the idea that modifying the Stokes drift formula by the inclusion of difference harmonics is a good idea, with the important caveat that all modes (including subharmonics) are assumed in deep water (see below).
Waves with more than two fundamental harmonics at lowest order become increasingly complex - from an algebraic perspective when deriving an analytical theory, and from a practical perspective when attempting to meaningfully describe the various parameter regimes. As with bichromatic waves, for focused wave groups we see strong localised displacements at depth in the opposite of the wave propagation direction, but find that the averaged drift remains positive, with slightly larger values for nonlinear theories that include the effects of subharmonics. We explore the consequences of extending the ubiquitous formulation of Kenyon (1969) by including difference harmonic terms for energy spectra, and obtain significant changes to surface Stokes drift and Stokes transport for several parametric spectra. These changes are strongly dependent on resolving high frequencies (see Lenain & Pizzo (2020)), and it would be interesting to see if such effects can be seen in field experiments or suitable ensemble-averaged flume experiments.
Our objective has been to provide a systematic and transparent exposition of drift associated with waves in nonlinear potential flow. To this end numerous simplifications have been made. Noteworthy among these is the neglect of viscous effects at the free surface, which has been known since work of Longuet-Higgins (1953) to play an important role. We have likewise neglected directional effects. Although our methodology can be adapted without change to include wave directional spreading, the additional freedom that this provides makes for a significant additional burden. Nevertheless, directional spreading is key when comparing with real sea states, and it would be of interest to follow in the footsteps of the recent work by Higgins et al. (2020) and explore these effects. Finally, and significantly, we have relied on the simplification of infinite water depth. This has allowed us to present numerous simple and compact formulations which are difficult to find in the existing literature, and contributed – we hope – to a clear presentation. Nevertheless, it is important to note that long bound harmonics may “feel the bottom” even when the free waves do not, and so the inclusion of finite depth is of central importance. Indeed, finite depth effects can also notably shape the particle drift for monochromatic waves, a subject which has been the subject of exploration from the work of Longuet-Higgins (1953) to the present day (Grue & Kolaas, 2017). We intend to treat finite depth waves in a subsequent work.
There are also numerous other cases of interest that should be mentioned. It might be argued that, after Stokes waves the next most-complex – and already unsteady – solution is that of standing waves. While some higher-order solutions are available in both Eulerian (Bryant & Stiassnie, 1994; Schwartz & Whitney, 1981) and Lagrangian (Chen & Hsu, 2009) variables, the theory is not nearly as complete as that for periodic, progressive waves. A systematic exploration of the kinematics of standing waves at various orders, in line with that suggested here for progressive waves, would be an interesting area for future work. Because of our assumptions of infinite water depth for all modes and a vanishing Eulerian mean flow it is quite difficult to compare with experimental results obtained in closed wave flumes, wherein Eulerian mean flows are typically depth-dependent, and therefore not irrotational. However, our methodology could be adapted to include uniform currents, and so enable a comparison of higher-order Eulerian and Lagrangian kinematics (the latter derived by Chen et al. (2012)) in this setting.
The fact that a weakly nonlinear theories in Eulerian variables generally make use of Taylor expansion to transfer the boundary conditions to a half-space (and therefore confine the fluid domain to ) has unfortunate consequences. It has given rise to a significant literature aimed to deal with the problem of evaluating kinematics in the surface zone between the crest and trough level, and particularly at wave crests, beginning with the stretching methods of Wheeler (1970) and continuing to the present day. Recent comparisons of these methods can be found, for example, in Stansberg et al (2008) or Johannessen (2010). We have not resolved the problem of identifying kinematics between crest and trough in Eulerian variables, but we have attempted a theoretically consistent, order-by-order exploration of a variety of waves.
With all of the caveats above, it is important to mention that the drift incorporated into operational models such as WAVEWATCH III is the simplest infinite-depth Stokes drift formulation (Li et al., 2016). This points to a continued need to better understand this formulation – with all the simplifications it involves – and how it may be further refined. We hope that the present work presents a modest step in this direction.
Appendix A On the non-uniqueness of Stokes expansion
The perturbation expansion associated with Stokes waves is often presented in textbooks as following a set sequence of steps. In fact, numerous choices are made, including whether to add linear terms in or to the potential, and how to treat constants appearing in the expansion. A useful discussion of some of these choices is found in (Svendsen, 2005, Ch. 8). We remark here on one aspect of this non-uniqueness that has received rather less attention (though see the comment in (Janssen, 2009, Appendix A)).
For a chosen linear wave amplitude there is a non-uniqueness in the Stokes’ expansion, which makes it necessary to specify the potential explicitly. For example, the following pairs of third-order solutions have the same linear amplitude
and
The former (44a)–(44b) is given by (Wehausen & Laitone, 1960, (27.25)), and appears naturally when the third order potential is set to zero. However, at each order solutions to the Laplace equation permit the addition of first-harmonic terms of the form . The latter (45a)–(45b) is the natural form obtained from the Zakharov formulation (see also (Janssen, 2009, (A.16)), (Gao et al., 2021, (6.7c))). Note that this flexibility in the perturbation expansion explains the seeming discrepancy identified by Zhang & Chen (1999) when matching these solutions.
When dealing with monochromatic waves, it is important to note that the solutions (44a)–(44b) and (45a)–(45b) need to be adjusted when a wave of fixed height is considered. According to these formulations, the height of the wave, defined as is related to the linear amplitude by either
where subscript 1 refers to (44a)–(44b) and subscript 2 to (45a)–(45b). For a prescribed height these expressions can be iteratively inverted to yield
Inserting this into the third order solution recovers the coefficients in terms of found, for example, in Fenton Fenton (1990).
Appendix B Coefficients of the higher-order potential
We provide here the coefficients of the second-order and third-order contributions to the potential from Section 2.3. The kernels appearing therein follow the usage of Krasitskii (1994).
Second-order coefficients are:
The third-order coefficients are:
As previously the shorthand notation is used to denote , and similarly for
Appendix C Effects of amplitude evolution on particle kinematics
In Section 2.2 we contend that it is appropriate to neglect the evolution of Fourier amplitudes while calculating particle trajectories and associated drift. This is based on the scale separation between the slow amplitude evolution (taking place on a time-scale of ) and the much faster particle motions, whose time-scale is in principle , i.e. proportional to the wave period. As we have seen, the motion of an individual particle depends on the wave phase, and so a single estimate of Stokes drift is – in some sense – an average value only. This is borne out by our methodology for calculating the horizontal drift of particles using 1st, 2nd, and 3rd order theory for monochromatic waves (Figure 4), bichromatic waves (Figure 8) and multichromatic waves (Figure 13).
We explore the role of amplitude evolution for an extreme example in Figure 17, which considers the evolution of a modulationally unstable wave with 1/m and (linear) steepness together with two initially small disturbances and for For clarity, our comparison excludes 2nd and 3rd order bound modes (see Sec. 2.3), and focuses only on the effects of amplitude evolution.
At the initial time the free surface is essentially monochromatic (Fig. 17(b)) and nearly all of the energy is the carrier wave . Subsequently, instability gives rise to energy exchange between the carrier and the side-bands and which can be captured by solving the Zakharov equation (8), and is shown in Figure 17(a). Note that the characteristic time for energy exchange, s, where is the carrier period.
Panel (c) of Figure 17 shows a particle trajectory at an initial depth of m. The particle is tracked for 10 starting at time s – indicated by black curves on the left of panel (a). The amplitude evolution here is such that the energy is being transferred from the carrier to the side-bands, until equipartition is reached in the Fourier amplitudes. The energy transfer is captured in Panel (d), which shows the initial Fourier amplitudes in red (carrier larger than side bands) and the final Fourier amplitudes after 10 in blue. In Panel (c) we compute the particle trajectory using both the initial amplitudes only (red curves) and with time-dependent amplitude evolution of Panel (a) (blue curves). Panel (e) shows the corresponding evolution away from equipartition in amplitudes, i.e. starting at as shown in Panel (f), and following the amplitude evolution for 10 along the black curves in Panel (a).
As alluded to earlier, this is an extreme case characterised by a very narrow spectrum and a dramatic spectral evolution from monochromatic waves to a strongly modulated wave train (see Panel (b)). Nevertheless, the amplitude evolution is essentially negligible over a single carrier period , and not dramatic even over time-scales of Moreover, any change in particle paths due to amplitude evolution depends both on the phase-position of the particle as well as where we are in the cycle of energy exchange.
In Figure 18 we consider a rather milder type of nonlinear interaction (though still significant enough to be interesting), where a Gaussian spectrum (with 1/m, 1/m, and m as per (40)) consisting of 10 Fourier modes with maximum initial slope is allowed to evolve according to the cubic Zakharov equation (8). The energy exchange between modes leads to changes in the packet-structure of the free surface (shown in Panel (a)), owing to the spectral broadening which is illustrated in Panel (e), which is solved for approximately .
Panels (b) and (c) of Figure 18 show particle paths evolving for 10 peak periods starting at and s, respectively. As in Figure 17, blue curves include the amplitude evolution captured by the Zakharov equation while red curves employ constant, initial amplitudes. Unsurprisingly, given the slow rate of spectral change, close to s there is scarce difference in the particle paths over 10 in panel (b). On the other hand, further into the time evolution the wave fields are clearly different, and particle paths cannot be expected to agree. Despite this, when averaging the drift of particle trajectories over 20 peak wavelengths we find the Stokes drift for evolving amplitudes agrees very well with the result for constant amplitudes, shown in Panel (d). In fact, due to the exclusion of bound modes we find that the Stokes drift is very close to the linear formula (I) from (41) – shown in yellow circles in panel (d).
Recent work by Blaser et al. (2025) studied the Stokes drift of wave packets undergoing both dispersive focusing and amplitude evolution using a fully nonlinear potential flow solver. Our focus in this appendix is somewhat different from theirs, as we do not concentrate on an isolated wave packet, nor examine the transport for the entire packet; nevertheless, the extent to which a simple cubically nonlinear theory such as that developed here could explain some of the findings from their numerical simulations remains an interesting question for future work.